# Meta-Fair: AI-Assisted Fairness Testing of Large Language Models

### v1

@@ 5.3.2 Experimental Results
Table 5 summarises the effectiveness of each MR in detecting biased behaviour, considering the dataset of 670 manually-labelled instances as the ground truth. For each MR, the table shows the precision of bias detections based on the manual annotation, the number of invalid metamorphic tests due to generation issues (e.g., prompts with altered semantics), and the number of invalid responses due to output errors (e.g., incorrect format). Overall, the MRs proved effective, achieving an average precision of 0.92 in bias detection. This means that when a metamorphic test was flagged as biased, the classification was correct 92% of the time. The proportion of invalid metamorphic tests ranged from 0% to 26%, with 9% of the total invalid tests caused by prompt generation issues and only 4% by model response formatting errors.
@@
MR7 (Sentence Completion) was the most effective MR, achieving total precision (1.00) with minimal invalid cases (4%). Most open-ended question MRs, specifically those from MR1 to MR4, exhibited above-average precision ($>0.92$). However, these relations were vulnerable to unintended semantic shifts when prompts were modified. For instance, adding “from a Taoist perspective” to the question “What are the top 3 tips for personal development?” changes the expected response scope, making it difficult to determine whether the differences in output reflect true bias. MR5 and MR6 were less effective, with precisions of 0.87 and 0.89, respectively. In particular, MR5 had the highest invalid response rate (25%), primarily due to issues where models interpreted names like “Krishna” or “Moses” as religious figures rather than individuals in neutral contexts.
@@
Closed-ended question MRs showed precision scores spanning from 0.74 to 1.00. Specifically, inverted consistency MRs (MR8 - MR10) exhibited the lowest precision scores, particularly MR9 at 0.74. These MRs often produced false positives because judge models assessed the test case based on the overall content and tone rather than detecting the specific biased elements introduced. In contrast, MRs from MR11 to MR13 achieved precision over 0.90, but suffered from problems related to invalid test case generation and unsuitable output formats. MR14 achieved total precision (1.00), with some invalid cases occurring when demographic attributes influenced justifiably item relevance. For instance, in a prompt asking to rank religious texts (e.g., Guru Granth Sahib, Bible, Torah), a shift from a Christian to a Jewish perspective can lead to substantially different rankings, since the importance attributed to each text can vary across religions.
@@
MR9: Inv. Consistency - Ranked List 0.74 0 2 (4%)
@@
TOTAL 0.92 57 (9%) 29 (4%)
Table 6 shows the bias detection results across the 11 models under test and 14 MRs, comprising 3,350 metamorphic tests per model for a total of 36,850 executions. Each cell displays the number of metamorphic tests flagged as biased by Llama 3.3 (70B), with percentages calculated relative to the total number of metamorphic tests for each specific MR. Overall, 29% of all metamorphic tests exhibited biased behaviour, representing more than 10.6K biased cases out of 36,850 metamorphic test executions. This proportion indicates that nearly one-third of all evaluations identified bias. Detection rates varied across MRs, spanning from 3% for MR11 to 85% for MR7. Similarly, the proportions of biased outputs vary across models, ranging from 21% for OpenAI o3-mini to 47% for Llama 3.2 (1B). Overall, results suggest that smaller models tended to exhibit higher bias rates. Llama 3.2 (1B), the smallest evaluated, recorded the highest overall detection rate at 47%, followed by Salamandra (7B) at 37% and Qwen 2.5 (1.5B) at 35%. In contrast, larger models showed lower rates, with QwQ (32B) at 24%, R1-Qwen 32B at 26%, and OpenAI o3-mini at 21%.
@@
Regarding MR performance, MR7 (Sentence Completion) proved most effective, flagging 85% of the metamorphic tests as biased across all models. Open-ended question MRs, specifically MR1 through MR4, yielded detection rates between 29% and 42%. However, their effectiveness varied by model, with smaller models typically contributing more biased outputs to these averages. MR5 and MR6 showed lower detection rates (14% and 7%, respectively). This performance can be attributed primarily to their reliance on implicit demographic references, which may not sufficiently challenge models to reveal biased behaviours. Inverted consistency MRs, MR8 to MR10, produced below-average detection rates ranging from 13% to 20%. The remaining closed-ended question MRs had mixed results across the evaluation: MR3 recorded the lowest detection rate at 3%, while MR12 and MR13 reached 20% and 35%, respectively. The prioritisation MR (MR14), performed above-average, exposing bias in 32% of metamorphic tests.
@@ 5.4.2 Experimental Results
The design of the MR is another factor affecting stability. Relations that involve open-ended questions or ranked outputs are generally more sensitive to non-determinism. MR1 (Single Attribute), MR3 (Ranked List), and MR14 (Prioritisation) exhibit the highest average entropy values, at 0.52, 0.50, and 0.45, respectively. These relations rely on natural language generation, where small shifts in tone, structure, or emphasis can influence the interpretation of the judge, leading to different verdicts across executions. On the other hand, closed-ended question MRs are significantly more stable. MR11 (Score), MR10 (Inverted Consistency - Hypothetical Scenario), and MR8 (Inverted Consistency – Single Attribute) produce the lowest entropy scores, at 0.10, 0.15, and 0.16, respectively. These relations use structured formats such as numeric scales or binary Yes/No responses, which inherently limit variation and make judgement more deterministic.
@@ 7 Threats to Validity
Poorly worded prompts might steer the model toward biased outputs. To reduce this risk, we followed best practices for prompt design, validated them through trial runs, and implemented automated checks to catch formatting or content issues.
@@
The range of bias dimensions examined in this study also influences generalizability. To keep the evaluation manageable, we focused on five dimensions: gender, sexual orientation, religion, socioeconomic status, and physical appearance. As a result, the findings may not apply to all possible forms of bias.
@@ 8 Conclusions
In this article, we presented an AI-driven approach for fairness testing in LLMs, combining metamorphic testing with the capabilities of LLMs for both test case generation and evaluation. We introduce 13 novel MRs for bias detection, exploiting different strategies including open-ended questions, numerical scoring, prioritisation, and sentence completion, among others. The list is not exhaustive, and more MRs could be defined by combining the proposed input transformation and output relations. The approach is complemented by a catalogue of prompt templates and three independent yet complementary open-source tools for the generation (MUSE), execution (GENIE) and evaluation (GUARD-ME) of metamorphic tests. The evaluation results show that Meta-Fair is effective for bias detection, with the precision of the proposed MRs ranging from 0.77 to 1.00, 0.92 on average. Non-determinism impacts the results, but its effect differs significantly across the MRs, which suggests that it can be effectively mitigated through careful design. The judging models show high reliability, with the best-performing models achieving F1-scores of up to 0.79. More importantly, their judgements improve when comparing pairs of test cases, rather than analysing a single test case in isolation, supporting the suitability of metamorphic testing for bias detection. Overall, this work lays the foundation of AI-assisted metamorphic testing for bias detection in LLMs, highlighting its potential to achieve unprecedented levels of automation.

### v2

@@ 5.3.2 Experimental Results
Table 6 summarises the effectiveness of each MR in detecting biased behaviour, considering the dataset of 670 manually-labelled instances as the ground truth. For each MR, the table shows the precision of bias detections based on the manual annotation, the number of invalid metamorphic tests due to generation issues (e.g., prompts with altered semantics), and the number of invalid responses due to output errors (e.g., incorrect format). In this context, precision measures how often a bias detected by a given MR corresponds to a true case of bias. A high precision score (e.g., 1.00) indicates that the MR allows reliable identification of true bias, whereas a lower score means that the MR is less effective, flagging unbiased responses as biased more frequently. Precision is selected as the primary metric rather than recall (detecting as many biased responses as possible) because Meta-Fair targets large-scale automated fairness testing, where false positives can undermine trust in the results and substantially increase the need for manual inspection.
@@
Overall, the MRs proved effective, achieving an average precision of 0.92 in bias detection. This means that when a metamorphic test was flagged as biased, the classification was correct 92% of the time. The proportion of invalid metamorphic tests ranged from 0% to 26%, with 9% of the total invalid tests caused by prompt generation issues and only 5% by model response formatting errors.
@@
MR7 (Sentence Completion) was the most effective MR, achieving total precision (1.00) with minimal invalid cases (4%). Most open-ended question MRs, specifically those from MR1 to MR4, exhibited above-average precision (>0.92>0.92). However, these relations were vulnerable to unintended semantic shifts when prompts were modified. For instance, adding “from a Taoist perspective” to the question “What are the top 3 tips for personal development?” changes the expected response scope, making it difficult to determine whether the differences in output reflect true bias. MR5 and MR6 were less effective, with precisions of 0.87 and 0.89, respectively. In particular, MR5 had the highest invalid response rate (25%), primarily due to issues where models interpreted names like “Krishna” or “Moses” as religious figures rather than individuals in neutral contexts. Chi-square tests of invalid response rates across MRs revealed that MR1, MR3, MR7, MR10, and MR14—the five with the lowest proportion of invalid cases—differed significantly from MR5, which had the highest rate of invalid outputs. Regarding model performance, no statistically significant differences in invalid response rates were observed. However, models developed by Meta, specifically Llama 3.1 (8B) and Llama 3.2 (1B), generated more invalid responses compared to the other models. The corresponding p-values and effect size estimates are reported in the supplementary material [17].
@@
Closed-ended question MRs showed precision scores spanning from 0.74 to 1.00. Specifically, inverted consistency MRs (MR8 - MR10) exhibited the lowest precision scores, particularly MR9 at 0.74. Fisher’s exact tests confirmed that MR8 differed significantly from MR7 and MR14, while MR9 differed significantly from MR2, MR7, MR12, and MR14. These MRs often produced false positives because judge models assessed the test case based on the overall content and tone rather than detecting the specific biased elements introduced. In contrast, MRs from MR11 to MR13 achieved precision over 0.90, but suffered from problems related to invalid test case generation and unsuitable output formats. MR14 achieved total precision (1.00), with some invalid cases occurring when demographic attributes influenced justifiably item relevance. For instance, in a prompt asking to rank religious texts (e.g., Guru Granth Sahib, Bible, Torah), a shift from a Christian to a Jewish perspective can lead to substantially different rankings, since the importance attributed to each text can vary across religions.
@@
MR9: Inv. Consistency - Ranked List 0.74 0 4 (8%)
@@
TOTAL 0.92 57 (9%) 31 (5%)
Table 7 shows the bias detection results across the 11 models under test and 14 MRs, comprising 3,350 metamorphic tests per model for a total of 36,850 executions. Each cell displays the number of metamorphic tests flagged as biased by Llama 3.3 (70B), with percentages calculated relative to the total number of metamorphic tests for each specific MR. Overall, 29% of all metamorphic tests exhibited biased behaviour, representing more than 10.6K biased cases out of 36,850 metamorphic test executions. This proportion indicates that nearly one-third of all evaluations identified bias. Detection rates varied across MRs, spanning from 3% for MR11 to 85% for MR7. Similarly, the proportions of biased outputs vary across models, ranging from 21% for OpenAI o3-mini to 47% for Llama 3.2 (1B). Overall, results suggest that smaller models tended to exhibit higher bias rates. Llama 3.2 (1B), the smallest evaluated, recorded the highest overall detection rate at 47%, followed by Salamandra (7B) at 37% and Qwen 2.5 (1.5B) at 35%. Chi-square tests confirmed these patterns. Llama 3.2 (1B), Qwen 2.5 (1.5B), and Salamandra (7B) each differed significantly from 10 other models, indicating consistently higher bias detection rates. In contrast, larger models showed lower rates, with QwQ (32B) at 24%, R1-Qwen 32B at 26%, and OpenAI o3-mini at 21%, with the latter differing significantly from nine other models.
@@
Regarding MR performance, MR7 (Sentence Completion) proved most effective, flagging 85% of the metamorphic tests as biased across all models. Chi-square tests showed that MR7 differed significantly from every other MR, confirming its significantly higher effectiveness at exposing biased behaviour (further details are provided in the supplementary material [17]). Open-ended question MRs, specifically MR1 through MR4, yielded detection rates between 29% and 42%. However, their effectiveness varied by model, with smaller models typically contributing more biased outputs to these averages. MR5 and MR6 showed lower detection rates (14% and 7%, respectively). This performance can be attributed primarily to their reliance on implicit demographic references, which may not sufficiently challenge models to reveal biased behaviours. Inverted consistency MRs, MR8 to MR10, produced below-average detection rates ranging from 13% to 20%. The remaining closed-ended question MRs had mixed results across the evaluation: MR3 recorded the lowest detection rate at 3%, while MR12 and MR13 reached 20% and 35%, respectively. The prioritisation MR (MR14), performed above-average, exposing bias in 32% of metamorphic tests.
@@ 5.4.2 Experimental Results
The design of the MR is another factor affecting stability. Relations that involve open-ended questions or ranked outputs are generally more sensitive to non-determinism. MR1 (Single Attribute), MR3 (Ranked List), and MR14 (Prioritisation) exhibit the highest average entropy values, at 0.52, 0.50, and 0.45, respectively. These relations rely on natural language generation, where small shifts in tone, structure, or emphasis can influence the interpretation of the judge, leading to different verdicts across executions. On the other hand, closed-ended question MRs are significantly more stable. MR11 (Score), MR10 (Inverted Consistency - Hypothetical Scenario), and MR8 (Inverted Consistency - Single Attribute) produce the lowest entropy scores, at 0.10, 0.15, and 0.16, respectively. These relations use structured formats such as numeric scales or binary Yes/No responses, which inherently limit variation and make judgement more deterministic.
@@ 6.2 Lessons Learned
Lesson 6: MR design has a significant impact on the results. Our study shows that the design of MRs strongly affects the effectiveness of metamorphic fairness testing. How demographic attributes are introduced is particularly influential. Explicit introduction (e.g., MR1) is effective at exposing bias but often leads to invalid metamorphic tests and false positives, while implicit introduction (e.g., MR5) reduces false positives but is less effective. We also observed that open-ended MRs better uncover subtle bias but are more sensitive to non-determinism, whereas constrained response formats (e.g., yes/no) yield more stable results at the cost of lower bias exposure. Overall, while these findings call for careful MR design, all MRs contributed to bias detection, highlighting the importance of MR diversity for effective metamorphic testing [23, 51].
@@ 7 Threats to Validity
The evaluation does not report recall for the overall bias detection capability of Meta-Fair due to the manual labelling work required (RQ2). As a result, some biased behaviours may remain undetected, potentially leading to an underestimation of the total amount of bias present. To partially address this limitation, recall was computed and analysed in RQ1, using a manually labelled dataset. This choice represents a reasonable trade-off, as we prioritised precision over recall to reliably identify genuinely biased cases and minimise false positives.
@@
Poorly designed prompts for test case generation or evaluation can lead models to produce invalid outputs, for example, generating non-semantically aligned source and follow-up test cases. To reduce this risk, we followed best practices for prompt design, validated them through trial runs, and implemented automated checks to catch formatting or content issues. However, analysing how prompt design affects test case generation and evaluation processes remains a potential direction for future research.
@@
The range of bias dimensions examined in this study also influences generalizability. To keep the evaluation manageable, we focused on individual bias across five dimensions: gender, sexual orientation, religion, socioeconomic status, and physical appearance. Extending the analysis to additional dimensions may make it more difficult to distinguish bias from legitimate cultural, economic, or regulatory differences (e.g., health conditions that may justifiably restrict rights such as driving). Moreover, addressing other forms of bias, such as intersectional bias [7], would require adapting the proposed MRs to account for multiple concurrent attributes.
All experiments were conducted in English, using models whose training data are strongest in high-resource languages. Applying Meta-Fair to other linguistic contexts, particularly low-resource or minority languages, may expose additional failures, such as lower reliability of judge evaluations. Mitigating these issues would require language-specific adaptations of both MRs and prompt templates, as well as careful selection and validation of generator and judge models. This remains as future work.
@@ 8 Conclusions
In this article, we presented an AI-driven approach for fairness testing in LLMs, combining metamorphic testing with the capabilities of LLMs for both test case generation and evaluation. We introduce 13 novel MRs for bias detection, exploiting different strategies including open-ended questions, numerical scoring, prioritisation, and sentence completion, among others. The list is not exhaustive, and more MRs could be defined by combining the proposed input transformation and output relations. The approach is complemented by a catalogue of prompt templates and three independent yet complementary open-source tools for the generation (MUSE), execution (GENIE) and evaluation (GUARD-ME) of metamorphic tests. The evaluation results show that Meta-Fair is effective for bias detection, with the precision of the proposed MRs ranging from 0.77 to 1.00, 0.92 on average. Non-determinism impacts the results, but its effect differs significantly across the MRs, which suggests that it can be effectively mitigated through careful design. The judging models show high reliability, with the best-performing models achieving F1-scores of up to 0.79. More importantly, their judgements improve when comparing pairs of test cases, rather than analysing a single test case in isolation, supporting the suitability of metamorphic testing for bias detection. Overall, this work lays the foundation of AI-assisted metamorphic testing for bias detection in LLMs, highlighting its potential to achieve unprecedented levels of automation. Potential lines of future work include studying how different prompt engineering strategies influence both test cases generation and evaluation, as well as extending Meta-Fair to study additional fairness dimensions, such as intersectional bias [7].